3 ms·
Quick trip to Wikipedia gives the reason I assume is correct, which is that we only notice Paul because of his success. After starting to notice him, the odds o
by edanm 16y ago
Quick trip to Wikipedia gives the reason I assume is correct, which is that we only notice Paul because of his success. After starting to notice him, the odds of guessing the remaining games was much more likely (assuming no cheating of any kind, even).
Quotes from the article (http://en.wikipedia.org/wiki/Paul_the_octopus http://en.wikipedia.org/wiki/Paul_the_octopus):
'Spiegelhalter and Roquain point out that there are "other animals that have attempted but failed to predict the outcome of football matches"; it is not remarkable that one animal is more successful than the others (including humans), and only the successful animals gain public attention after the fact'
'Paul started to receive international recognition after he correctly predicted Germany's win over England; after that he made four correct predictions. The chance of those final four predictions being correct is 1/24 = 6.25% (odds 15 to 1).'